OpenStudy (anonymous):

Your piggy bank has a total of 47 coins in it; some are dimes and some are nickels. If you have a total of $3.95, how many nickels do you have?

OpenStudy (anonymous):

there are many answers...

OpenStudy (anonymous):

Wait...no

OpenStudy (amistre64):

n = (3.95 - 47(10))/(5-10) perhaps

OpenStudy (anonymous):

Hold on, brb

OpenStudy (amistre64):

that almost works :)

OpenStudy (amistre64):

.10 and .05

OpenStudy (amistre64):

n = (3.95 - 47(.10))/(.05-.10) is better

OpenStudy (amistre64):

15 is good

OpenStudy (anonymous):

make 2 equations n+d=47 and 0.05n+0.10d=3.95

OpenStudy (anonymous):

It's 32 dimes and 15 nickels, my bad